# Mathematical Language Models: A Survey

### v3

@@ 4.4. Advanced CoT Methods
For LLMs, research shows that the Socratic teaching method can also improve their abilities to a certain extent. For example, the framework of Socratic Models (SMs) (Zeng et al., 2022) significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. In Chang et al. (Chang, 2023), techniques including definition, dialectics, cross-examination, induction, and counterfactual reasoning are utilized. These techniques contribute to enhancing the LLMs’ comprehension of user intentions, thereby improving model’s effectiveness in addressing user queries. The inductive, deductive, and analogical reasoning techniques in the Socratic teaching method establish a comprehensive analytical framework for solving mathematical problems (Chang, 2023; Al-Hossami et al., 2023a), ensuring a strong and concise ability to address the complexities of mathematical problem-solving. Moreover, contributions from Ang et al. (Ang et al., 2023) and Al et al. (Al-Hossami et al., 2023b) provided invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability.
Overall, the integration of Socratic teaching techniques in LLMs offers novel perspectives and tools for mathematics education. Unlike CoT (Wei et al., 2022), Socratic questioning explicitly guides the thinking process, stimulating effective recursive thinking and displaying greater resilience to errors in the thinking process (Qi et al., 2023). Ranging from basic conversational guidance to intricate problem-solving, these methods demonstrate the effective fusion of modern technology with classical ideas to bolster the learning and comprehension abilities of LLMs.
@@
VERBPHYSICS (Forbes and Choi, 2017) 733 1,096 1,828 3,657 EN Calculation Training Formula
Clinical (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) 11,170 1,625 3,220 16,015 EN Calculation Training Formula
Scientific (Spithourakis and Riedel, 2018) 14,694 2,037 4,231 20,962 EN Calculation Training Formula
DoQ (Elazar et al., 2019) 587 5,418 6,007 12,012 EN Calculation Training Formula
DROP (Dua et al., 2019) 77,409 9,536 9,622 96,567 EN Calculation Training Text
AddSub (Hosseini et al., 2014) - - - 395 EN MWP Training Formula
SingleOp (Roy et al., 2015) 265 107 159 531 EN MWP Training Formula
SingleEq (Koncel-Kedziorski et al., 2015) - - - 508 EN MWP Training Formula
MultiArith (Roy and Roth, 2015) 420 - 180 600 EN MWP Training Formula
Alg514 (Kushman et al., 2014) - - - 514 EN MWP Training Formula
Math23k (Wang et al., 2017) 22,162 - 1000 23,162 CH MWP Training Formula
AQuA (Ling et al., 2017) 97,467 254 254 97,975 EN MWP Training Text
MathQA (Amini et al., 2019) 29,837 4,475 28,985 37,297 EN MWP Training Formula
GSM8K (Cobbe et al., 2021) 7,473 1,319 8,792 EN MWP Training Text
SVAMP (Patel et al., 2021) 700 300 1,000 EN MWP Training Formula
DRAW (Upadhyay and Chang, 2015) - - - 1,000 EN MWP Training Formula
Dolphin1878 (Shi et al., 2015) - 374 1,504 1,878 EN MWP Training Formula
AllArith (Roy and Roth, 2017) - - - 831 EN MWP Training Formula
DRAW-1K (Upadhyay and Chang, 2017) - - - 1,000 EN MWP Training Formula
HMWP (Qin et al., 2020) - - - 5,470 CH MWP Training Formula
ArMATH (Alghamdi et al., 2022) - - - 6,000 Arabic MWP Training Formula
TabMWP (Lu et al., 2023b) - - - 38,431 EN MWP Training Text
Ape210K (Zhao et al., 2020) - - - 210,488 CH MWP Training Formula
TAL-SCQ5K-CH  [^1] https://github.com/math-eval/TAL-SCQ5K 3,000 - 2,000 5,000 CH MWP Training Text
TAL-SCQ5K-EN  [^2] https://github.com/math-eval/TAL-SCQ5K 3,000 - 2,000 5,000 EN MWP Training Text
FinQA (Chen et al., 2021) 6,251 883 1,147 8,281 EN MWP Training Formula
REALFP (Kalyan et al., 2021) 185 185 558 928 EN MWP Training Formula
SYNTHFP (Kalyan et al., 2021) 10,000 1,000 1,000 12,000 EN MWP Training Formula
TAT-QA (Zhu et al., 2021) - - - 16,552 EN MWP Training Text
MultiHiertt (Zhao et al., 2022) 7,830 1,044 1,566 10,440 EN MWP Training Formula
MATHPILE (Wang et al., 2023c) - - - 903,180 EN MWP Training Text
OpenWebMath (Paster et al., 2023) - - - - EN MWP Training Formula
MML (Grabowski et al., 2015) - - - 57,882 EN TP Training Formula
HolStep (Kaliszyk et al., 2017) 2,013,046 - 196,030 2,209,076 EN TP Training Formula
Feit-Thompson (Huang et al., 2019) - - - 83,478 EN TP Training Formula
CoqGym (Yang and Deng, 2019) - - - 71,000 EN TP Training Formula
HOList (Bansal et al., 2019) - - - 29,462 EN TP Training Formula
IsarStep (Li et al., 2021) 820,000 5,000 5,000 830,000 EN TP Training Formula
LISA (Jiang et al., 2021) - - - 183,000 EN TP Training Formula
NaturalProofs (Welleck et al., 2021) 32,000 EN TP Training Text
LeanStep (Han et al., 2022) - - - 21,606,000 EN TP Training Formula
NumGLUE (Mishra et al., 2022b) - - - 101,835 EN Calculation Benchmark Text
Dophin18k (Huang et al., 2016) - - - 18,460 EN MWP Benchmark Text
MAWPS (Koncel-Kedziorski et al., 2016) - - - 3,320 EN MWP Benchmark Formula
ASDiv (Miao et al., 2020) - - - 2,305 EN MWP Benchmark Formula
MATH (Hendrycks et al., 2021b) 7,500 5,000 12,500 EN MWP Benchmark Text
MGSM (Shi et al., 2022) - - - - Multilingual MWP Benchmark Text
Mathematics (Saxton et al., 2019) 2,000,000 100,000 2,100,000 EN MWP Benchmark Formula
MMLU-Math (Hendrycks et al., 2021a) - - - 906 EN WMP Benchmark Formula
AGIEval (Zhong et al., 2023) - - - 469/220 CH/EN MWP Benchmark Formula
MATH 401 (Yuan et al., 2023b) - - - 401 EN MWP Benchmark Formula
INT (Wu et al., 2021) - - - - EN TP Benchmark Formula
miniF2F (Zheng et al., 2023) - 244 244 488 EN TP Benchmark Formula
Aggregate (Roy and Roth, 2018) - - - 1,492 EN MWP Augmented Formula
MathQA-Python (Austin et al., 2021) 19,209 2,822 2,822 23,914 EN MWP Augmented Code
Math50k (Li et al., 2023a) - - - 50,000 EN WMP Augmented Text
PRM800K (Lightman et al., 2023) - - - 2,868 EN WMP Augmented Text
MetaMathQA (Yu et al., 2023b) - - - 395,000 EN MWP Augmented Text
Lila (Mishra et al., 2022a) - - - 134,000 EN MWP Augmented Code
PEN (Kim et al., 2022) - - - 3,581 EN MWP Augmented Formula
miniF2F+informal (Jiang et al., 2023b) - 244 244 488 EN TP Augmented Formula
NaturalProofs-Gen (Welleck et al., 2022) 12,500 1,000 1,000 14,500 EN TP Augmented Text
@@
Table 1. The statistics information of mathematical datasets. Solution means the format of the output, such as text, formula and code.
@@ 5. Datasets
To train and evaluate the arithmetic and math ability of language models, various math word problems datasets (Saxton et al., 2019; Hendrycks et al., 2021b; Cobbe et al., 2021; Kushman et al., 2014) are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1).
@@ 5.1.1. Mathematical Calculation
Several studies have introduced datasets aiming to identify numerical information within text (Spithourakis et al., 2016; Spithourakis and Riedel, 2018; Elazar et al., 2019), specifically targeting the prediction of attribute values. The Clinical Data (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) consists of clinical records sourced from the London Chest Hospital. Each patient record includes a text report and structured KB tuples detailing 20 potential numeric attributes like age and gender. Scientific Data (Spithourakis and Riedel, 2018) encompasses paragraphs extracted from Cornell’s ARXIV repository, featuring over half a million converted papers across 37 scientific sub-fields. Additionally, Elazar et al. (Elazar et al., 2019) proposed the Distributions over Quantities (DoQ) dataset, comprising empirical counts of scalar attribute values linked to more than 350K nouns, adjectives, and verbs across 10 different attributes.
Furthermore, certain studies consider the relational knowledge between various numbers. For instance, the VERBPHYSICS (Forbes and Choi, 2017) dataset aggregates crowdsourced information about actions and objects, encompassing the relative knowledge of grounded object pairs and implications of actions associated with those objects. More intricately, the DROP (Dua et al., 2019) dataset necessitates Discrete Reasoning Over the content of Paragraphs. In this benchmark of 55,000 adversarially-created questions, a system is tasked with resolving references within a question, potentially spanning multiple input positions, and performing discrete operations like addition, counting, or sorting.
@@ 5.1.2. Math Word Problems
A large number of datasets are proposed for MWPs. Hosseini et al. (Hosseini et al., 2014) curated the AddSub dataset, primarily focusing on addition and subtraction problems (Figure 9). This dataset identifies relevant variables and their respective values to translate sentences into problem statements, represented as equations. Another dataset, SingleOp (Roy et al., 2015), encompasses a wider range of mathematical operations including multiplication and division. Its purpose is to facilitate reasoning about quantities articulated in natural language. DRAW (Upadhyay and Chang, 2015) comprises 1000 algebra word problems semi-automatically annotated for evaluating automatic solvers. It features gold coefficient alignments crucial for uniquely identifying equation system derivations. The Alg514 (Kushman et al., 2014) dataset comprises 514 linear algebra problems structured around 28 equation templates. Meanwhile, MultiArith (Roy and Roth, 2015) is designed to handle arithmetic problems involving multiple steps and operations, without predefined templates. MATHPILE (Wang et al., 2023c) is a diverse and high-quality math-centric corpus comprising about 9.5 billion tokens. OpenWebMath (Paster et al., 2023) is an open dataset inspired by these works containing 14.7B tokens of mathematical webpages from Common Crawl.
Several studies have introduced methods focused on predicting equations through semantic parsing. The SingleEq (Koncel-Kedziorski et al., 2015) dataset revolves around grade-school algebra word problems. It comprises samples that can be correlated to a single equation involving multiple mathematical operations on one variable, structured as a parsing tree (Figure 10). In contrast, the Dolphin1878 (Shi et al., 2015) dataset contains 1,878 word problems in mathematics, wherein each sample might encompass multiple equations and variables depicted through multiple trees. AllArith (Roy and Roth, 2017) offers a concise representation of dependencies among number units mentioned in a given problem, referred to as Unit Dependency Graphs (UDGs). To augment dataset size, Math23k (Wang et al., 2017) gathers 23,161 problems labeled with structured equations and corresponding answers. GSM8K (Cobbe et al., 2021) comprises 8.5K high-quality grade school math problems crafted by human problem writers. These problems typically involve 2 to 8 steps for resolution, predominantly requiring a sequence of elementary calculations using basic arithmetic operations to arrive at the final solution.
Figure 9. An example of AddSub, taken from (Hosseini et al., 2014).
Figure 10. An example of SingleEQ, taken from (Roy et al., 2015).
HMWP (Qin et al., 2020) aims to enhance dataset diversity by extracting math word problems from a Chinese K12 problem bank. This effort sought to validate the universality of math word problem solvers and expand research in MWPs to better align with real-world scenarios. The dataset encompasses three types of MWPs: arithmetic word problems, equation set problems, and non-linear equation problems. In total, it comprises 5,491 MWPs, including 2,955 single-variable linear MWPs, 1,636 two-variable linear MWPs, and 900 single-variable nonlinear MWPs. Additionally, MathQA (Amini et al., 2019) presents a diverse collection of 37,000 English multiple-choice math word problems spanning various mathematical domains (Figure 11).
Figure 11. An example of MathQA, taken from (Amini et al., 2019).
Existing research endeavors to enhance the comprehensibility and precision of intricate reasoning by providing fine-grained annotations of reasoning processes (Upadhyay and Chang, 2017; Ling et al., 2017; Chen et al., 2021; Lightman et al., 2023; Zhao et al., 2022). DRAW-1K (Upadhyay and Chang, 2017) introduces a novel dataset featuring 1,000 general algebra word problems, each annotated with derivations. AQUA (Ling et al., 2017) structures each question into four components: problem description, multiple-choice answer options, rationale, and correct option label (Figure 12). FinQA (Chen et al., 2021) presents an expert-annotated dataset comprising 8,281 financial QA pairs, meticulously annotated with numerical reasoning processes to ensure comprehensive explication. PRM800K (Lightman et al., 2023) assigns labels (positive, negative, or neutral) to each step in solving MATH problems sampled by a large-scale generator. The training set encompasses 800K step-level labels across 75K solutions to 12K problems.
Figure 12. An example of AQUA, taken from (Ling et al., 2017).
Additionally, Patel et al. (Patel et al., 2021) introduced the SVAMP challenge set to establish a more robust evaluation framework for methods designed to tackle elementary-level math word problems. Alghamdi et al. (Alghamdi et al., 2022) contributed the inaugural large-scale dataset, ArMATH, for Arabic MWPs. This dataset comprises 6,000 samples of primary-school math problems written in Modern Standard Arabic. Kalyan et al. (Kalyan et al., 2021) proposed two datasets, REALFP and SYNTHFP, for a novel task called Fermi Problems (FPs). FPs entail questions whose answers can only be approximately estimated due to the impracticality or impossibility of precise computation. This task necessitates the amalgamation of various reasoning abilities, including suitable abstractions, commonsense knowledge, and the creative synthesis of problem-solving strategies.
Recently, mathematical reasoning has incorporated multiple input modalities such as textual and tabular data (Lu et al., 2023b; Zhu et al., 2021; Zhao et al., 2022). For instance, TabMWP (Lu et al., 2023b) presents a new dataset comprising 38,431 open-domain grade-level problems that necessitate mathematical reasoning across both textual and tabular data. Zhu et al. (Zhu et al., 2021) introduced a large-scale QA dataset, TAT-QA, encompassing both tabular and textual data. This dataset often requires numerical reasoning—such as addition, subtraction, multiplication, division, counting, comparison/sorting, and their combinations—based on real financial reports to infer answers. Additionally, MultiHiertt (Zhao et al., 2022) is a proposed dataset consisting of QA pairs derived from Multi Hierarchical Tabular and Textual data (Figure 13). Each document in this dataset contains multiple tables and longer unstructured texts, accompanied by fine-grained annotations of reasoning processes and supporting facts to elucidate complex numerical reasoning.
Figure 13. An example of MultiHiertt, taken from (Zhao et al., 2022).

### v4

@@ 4.4. Advanced CoT Methods
The Socratic teaching method (Nelson, 1980), derived from the philosophies of ancient Greek thinker Socrates, provides a unique perspective when considering the integration of LMs in mathematics education (Davies et al., 2021). This method, emphasizing questioning and dialogue, aims to guide learners toward profound thinking (Davis and Steinglass, 1997; Brickhouse and Smith, 2009; Ding et al., 2024). In realm of mathematics education, employing this method can aid students in grasping concepts with clarity, fostering more rigorous logical application, and promoting reflection and self-discovery. For LLMs, research shows that the Socratic teaching method can also improve their abilities to a certain extent. For example, the framework of Socratic Models (SMs) (Zeng et al., 2022) significantly broadens the scope of interactive capabilities. SMs facilitate the amalgamation of multiple LLMs to tackle novel multi-modal tasks using structured Socratic dialogues. In Chang et al. (Chang, 2023), techniques including definition, dialectics, cross-examination, induction, and counterfactual reasoning are utilized. These techniques contribute to enhancing the LLMs’ comprehension of user intentions, thereby improving model’s effectiveness in addressing user queries. The inductive, deductive, and analogical reasoning techniques in the Socratic teaching method establish a comprehensive analytical framework for solving mathematical problems (Chang, 2023; Al-Hossami et al., 2023a), ensuring a strong and concise ability to address the complexities of mathematical problem-solving. MATHDIAL (Macina et al., 2023) enables LLMs to engage in multi-round dialogues as both teacher and student, creating a conversational dataset that incorporates rich educational socratic attributes and background in mathematical reasoning problems. Moreover, contributions from Ang et al. (Ang et al., 2023) and Al et al. (Al-Hossami et al., 2023b) provided invaluable datasets and evaluation methodologies for Socratic questioning. These contributions underscore the practical applications and potential impact of Socratic methods in counseling and coaching, showcasing their diverse applicability. Overall, the integration of Socratic teaching techniques in LLMs offers novel perspectives and tools for mathematics education. Unlike CoT (Wei et al., 2022), Socratic questioning explicitly guides the thinking process, stimulating effective recursive thinking and displaying greater resilience to errors in the thinking process (Qi et al., 2023). Ranging from basic conversational guidance to intricate problem-solving, these methods demonstrate the effective fusion of modern technology with classical ideas to bolster the learning and comprehension abilities of LLMs.
4.5. Multi-modal Methods
Building on LMs, some multimodal approaches incorporate visual encoders, enabling the models to handle mathematics problems that rely on visual elements, such as geometry, topology, tabular statistics, etc. GPT-4V (OpenAI, 2023) is the first successfully proposed Multimodal LLMs (MLLMs) that demonstrate strong visual reasoning capabilities to handle certain image-based mathematical problems. Subsequently, a lot of other high-performing MLLMs have been introduced, including Gemini (Team et al., 2023), Qwen2-VL (Wang et al., 2024a), CogVLM (Wang et al., 2023b), GPT-4o (Hurst et al., 2024), GLM-4V (GLM et al., 2024) and InternVL (Chen et al., 2024d), etc. UnAC (wang and Fu, 2024) introduces a multimodal prompting approach that decomposes the reasoning process into three stages: understanding the image, abstracting image information, and systematically verifying the reasoning steps. Math-LLaVA (Shi et al., 2024) introduces a high-quality, diverse multimodal mathematical dataset called MathV360K, which significantly improves the model’s multimodal mathematical reasoning capabilities. LLaVA-o1 (Xu et al., 2024b) and AtomThink (Xiang et al., 2024) introduce long CoT into the domain of MLLMs. M-STAR (Anonymous, 2024) investigates the application of self-evolving training in multimodal reasoning by combining continuous self-evolving training, reward model re-ranking, and adaptive exploration strategies.
Table 1. The statistics information of mathematical datasets. Level: EE = Elementary, MM = Middle School, HH = High School, UU = University, CC = Competition, HH = Hybrid, OO = Others. Modality: TT = Text, MM = Multimodal
@@ 4.5. Multi-modal Methods
VERBPHYSICS (Forbes and Choi, 2017) 733 1,096 1,828 3,657 EN Calculation Training Formula - TT
Clinical (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) 11,170 1,625 3,220 16,015 EN Calculation Training Formula OO TT
Scientific (Spithourakis and Riedel, 2018) 14,694 2,037 4,231 20,962 EN Calculation Training Formula OO TT
DoQ (Elazar et al., 2019) 587 5,418 6,007 12,012 EN Calculation Training Formula - TT
DROP (Dua et al., 2019) 77,409 9,536 9,622 96,567 EN Calculation Training Text EE TT
AddSub (Hosseini et al., 2014) - - - 395 EN MWP Training Formula EE TT
SingleOp (Roy et al., 2015) 265 107 159 531 EN MWP Training Formula EE TT
SingleEq (Koncel-Kedziorski et al., 2015) - - - 508 EN MWP Training Formula EE TT
MultiArith (Roy and Roth, 2015) 420 - 180 600 EN MWP Training Formula - TT
Alg514 (Kushman et al., 2014) - - - 514 EN MWP Training Formula - TT
Math23k (Wang et al., 2017) 22,162 - 1000 23,162 CH MWP Training Formula EE TT
AQuA (Ling et al., 2017) 97,467 254 254 97,975 EN MWP Training Text UU TT
GSM8K (Cobbe et al., 2021) 7,473 1,319 8,792 EN MWP Training Text EE TT
SVAMP (Patel et al., 2021) 700 300 1,000 EN MWP Training Formula EE TT
DRAW (Upadhyay and Chang, 2015) - - - 1,000 EN MWP Training Formula OO TT
Dolphin1878 (Shi et al., 2015) - 374 1,504 1,878 EN MWP Training Formula - TT
HMWP (Qin et al., 2020) - - - 5,470 CH MWP Training Formula - TT
ArMATH (Alghamdi et al., 2022) - - - 6,000 Arabic MWP Training Formula EE TT
TabMWP (Lu et al., 2023c) - - - 38,431 EN MWP Training Text EEMM TT
TAL-SCQ5K  [^1] https://github.com/math-eval/TAL-SCQ5K 3,000 - 2,000 5,000 CH/EN MWP Training Text CC TT
REALFP (Kalyan et al., 2021) 185 185 558 928 EN MWP Training Formula CC TT
SYNTHFP (Kalyan et al., 2021) 10,000 1,000 1,000 12,000 EN MWP Training Formula OO TT
MultiHiertt (Zhao et al., 2022) 7,830 1,044 1,566 10,440 EN MWP Training Formula OO MM
MATHPILE (Wang et al., 2023e) - - - 903,180 EN MWP Training Text HH TT
OpenWebMath (Paster et al., 2023) - - - - EN MWP Training Formula HH TT
MathQA (Amini et al., 2019) 29,837 4,475 28,985 37,297 EN MQA Training Formula - TT
FinQA (Chen et al., 2021a) 6,251 883 1,147 8,281 EN MQA Training Formula OO TT
TAT-QA (Zhu et al., 2021) - - - 16,552 EN MQA Training Text OO MM
MML (Grabowski et al., 2015) - - - 57,882 EN TP Training Formula - TT
HolStep (Kaliszyk et al., 2017) 2,013,046 - 196,030 2,209,076 EN TP Training Formula - TT
CoqGym (Yang and Deng, 2019) - - - 71,000 EN TP Training Formula - TT
HOList (Bansal et al., 2019) - - - 29,462 EN TP Training Formula - TT
IsarStep (Li et al., 2021) 820,000 5,000 5,000 830,000 EN TP Training Formula UU TT
LISA (Jiang et al., 2021) - - - 183,000 EN TP Training Formula - TT
NaturalProofs (Welleck et al., 2021) 32,000 EN TP Training Text OO TT
LeanStep (Han et al., 2022) - - - 21,606,000 EN TP Training Formula - TT
NumGLUE (Mishra et al., 2022b) - - - 101,835 EN Calculation Benchmark Text HH TT
Dophin18k (Huang et al., 2016) - - - 18,460 EN MWP Benchmark Text EE TT
MAWPS (Koncel-Kedziorski et al., 2016) - - - 3,320 EN MWP Benchmark Formula HH TT
ASDiv (Miao et al., 2020) - - - 2,305 EN MWP Benchmark Formula EE TT
MATH (Hendrycks et al., 2021b) 7,500 5,000 12,500 EN MWP Benchmark Text HHOO TT
MGSM (Shi et al., 2022) - - - - Multilingual MWP Benchmark Text EE TT
Mathematics (Saxton et al., 2019) 2,000,000 100,000 2,100,000 EN MWP Benchmark Formula - TT
MMLU-Math (Hendrycks et al., 2021a) - - - 906 EN WMP Benchmark Formula EEHHCC TT
AGIEval (Zhong et al., 2023) - - - 469/220 CH/EN MWP Benchmark Formula HHCCOO TT
AIME [^2] https://www.kaggle.com/datasets/hemishveeraboina/aime-problem-set-1983-2024 - - - 933 EN MWP Benchmark Formula CC TT
MATHTRAP (Zhao et al., 2024b) - - - 105/155 EN MWP Benchmark Formula HH TT
MathVista (Lu et al., 2023a) - - - 6,141 EN MWP Benchmark Formula EEMMHHUU MM
Math-V (Wang et al., 2024b) - - - 3,040 EN MWP Benchmark Formula EEHHOO MM
CMM-Math (Liu et al., 2024) 22,248 - 5,821 28,069 CH MWP Benchmark Text EEMMHH MM
MathVerse (Zhang et al., 2025) - - - 2,612 EN MWP Benchmark Formula HH MM
GeoQA (Chen et al., 2021b) 3,499 745 754 4,998 EN MWP Benchmark Formula MM MM
M3CoT (Chen et al., 2024b) - - - 11,459 EN MWP Benchmark Formula OO MM
MR-MATH (Xia et al., 2024) - - - - EN MWP Benchmark Formula EE TT
MMMU-Math (Yue et al., 2024) - - - 505 EN MWP Benchmark Text UU MM
We-Math (Qiao et al., 2024) - - - 1700 EN MWP Benchmark Formula EEHHUU MM
U-MATH (Chernyshev et al., 2024) - - - 1,080 EN MWP Benchmark Formula UU MM
INT (Wu et al., 2021) - - - - EN TP Benchmark Formula - TT
miniF2F (Zheng et al., 2023) - 244 244 488 EN TP Benchmark Formula EEHHOO TT
FOLIO (Han et al., 2024) - - - 1,430 EN TP Benchmark Formula HH TT
Aggregate (Roy and Roth, 2018) - - - 1,492 EN MWP Augmented Formula HH TT
MathQA-Python (Austin et al., 2021) 19,209 2,822 2,822 23,914 EN MWP Augmented Code - TT
Math50k (Li et al., 2023a) - - - 50,000 EN WMP Augmented Text - TT
PRM800K (Lightman et al., 2024) - - - 2,868 EN WMP Augmented Text HHOO TT
MetaMathQA (Yu et al., 2023b) - - - 395,000 EN MWP Augmented Text HH TT
Lila (Mishra et al., 2022a) - - - 134,000 EN MWP Augmented Code HH TT
PEN (Kim et al., 2022) - - - 3,581 EN MWP Augmented Formula HH TT
miniF2F+informal (Jiang et al., 2023b) - 244 244 488 EN TP Augmented Formula EEHHOO TT
NaturalProofs-Gen (Welleck et al., 2022) 12,500 1,000 1,000 14,500 EN TP Augmented Text OO TT
@@ 5. Datasets
To train and evaluate the arithmetic and math ability of language models, various math word problems (MWPs) datasets (Saxton et al., 2019; Hendrycks et al., 2021b; Cobbe et al., 2021; Kushman et al., 2014) are designed. In this paper, we organize the frequently used datasets for mathematical language models by dividing them into training, benchmark and data augmentation datasets (Table 1).
@@ 5.1.1. Mathematical Calculation
Several studies have introduced datasets aiming to identify numerical information within text (Spithourakis et al., 2016; Spithourakis and Riedel, 2018; Elazar et al., 2019), specifically targeting the prediction of attribute values. The Clinical Data (Spithourakis et al., 2016; Spithourakis and Riedel, 2018) consists of clinical records sourced from the London Chest Hospital. Each patient record includes a text report and structured KB tuples detailing 20 potential numeric attributes like age and gender. Scientific Data (Spithourakis and Riedel, 2018) encompasses paragraphs extracted from Cornell’s ARXIV repository, featuring over half a million converted papers across 37 scientific sub-fields. Additionally, Elazar et al. (Elazar et al., 2019) proposed the Distributions over Quantities (DoQ) dataset, comprising empirical counts of scalar attribute values linked to more than 350K nouns, adjectives, and verbs across 10 different attributes. Furthermore, certain studies consider the relational knowledge between various numbers. For instance, VERBPHYSICS (Forbes and Choi, 2017) aggregates crowdsourced data on actions and objects, including knowledge of grounded object pairs and actions related to them. More intricately, DROP (Dua et al., 2019) dataset necessitates Discrete Reasoning Over the content of Paragraphs. In this benchmark of 55,000 adversarial questions, a system resolves references within a question across multiple positions and performs operations like addition, counting, or sorting.
@@ 5.1.2. Math Word Problems
A large number of datasets are proposed for MWPs. Hosseini et al. (Hosseini et al., 2014) curated AddSub, primarily focusing on addition and subtraction problems. This dataset identifies relevant variables and their respective values to translate sentences into problem statements, represented as equations. SingleOp (Roy et al., 2015), encompasses a wider range of mathematical operations including multiplication and division. Its purpose is to facilitate reasoning about quantities articulated in natural language. DRAW (Upadhyay and Chang, 2015) comprises 1000 algebra word problems semi-automatically annotated for evaluating automatic solvers. It features gold coefficient alignments crucial for uniquely identifying equation system derivations. The Alg514 (Kushman et al., 2014) dataset comprises 514 linear algebra problems structured around 28 equation templates. Meanwhile, MultiArith (Roy and Roth, 2015) is designed to handle arithmetic problems involving multiple steps and operations, without predefined templates. MATHPILE (Wang et al., 2023e) is a diverse and high-quality math-centric corpus comprising about 9.5 billion tokens. OpenWebMath (Paster et al., 2023) is an open dataset inspired by these works containing 14.7B tokens of mathematical webpages from Common Crawl. Several studies have introduced methods focused on predicting equations through semantic parsing. SingleEq (Koncel-Kedziorski et al., 2015) revolves around grade-school algebra word problems. It comprises samples that can be correlated to a single equation involving multiple mathematical operations on one variable, structured as a parsing tree. In contrast, Dolphin1878 (Shi et al., 2015) contains 1,878 word problems in mathematics, wherein each sample might encompass multiple equations and variables depicted through multiple trees. AllArith (Roy and Roth, 2017) offers a concise representation of the dependencies among number units mentioned in a given problem, referred to as Unit Dependency Graphs (UDGs). To augment dataset size, Math23k (Wang et al., 2017) gathers 23,161 problems labeled with structured equations and corresponding answers. GSM8K (Cobbe et al., 2021) comprises 8.5K high-quality grade school math problems crafted by human problem writers, where the problems typically involve 2 to 8 steps for resolution. HMWP (Qin et al., 2020) aims to enhance dataset diversity by extracting MWPs from a Chinese K12 problem bank. This effort sought to validate the universality of math word problem solvers and expand research in MWPs to better align with real-world scenarios. In total, it comprises 5,491 MWPs, including 2,955 single-variable linear MWPs, 1,636 two-variable linear MWPs, and 900 single-variable nonlinear MWPs. Additionally, MathQA (Amini et al., 2019) presents a diverse collection of 37,000 English multiple-choice MWPs spanning various mathematical domains. Existing research endeavors to enhance the comprehensibility and precision of intricate reasoning by providing fine-grained annotations of reasoning processes (Upadhyay and Chang, 2017; Ling et al., 2017; Chen et al., 2021a; Lightman et al., 2024; Zhao et al., 2022). DRAW-1K (Upadhyay and Chang, 2017) introduces a novel dataset featuring 1,000 general algebra word problems, each annotated with derivations. AQUA (Ling et al., 2017) structures each question into four components: problem description, multiple-choice answer options, rationale, and correct option label. FinQA (Chen et al., 2021a) presents an expert-annotated dataset comprising 8,281 financial QA pairs, meticulously annotated with numerical reasoning processes to ensure comprehensive explication. PRM800K (Lightman et al., 2024) assigns labels (positive, negative, or neutral) to each step in solving MATH problems sampled by a large-scale generator. The training set encompasses 800K step-level labels across 75K solutions to 12K problems. Additionally, Patel et al. (Patel et al., 2021) introduced the SVAMP challenge set to establish a more robust evaluation framework for methods designed to tackle elementary-level MWPs. Alghamdi et al. (Alghamdi et al., 2022) contributed the inaugural large-scale dataset, ArMATH, for Arabic MWPs. This dataset comprises 6,000 samples of primary-school math problems written in Modern Standard Arabic. Kalyan et al. (Kalyan et al., 2021) proposed two datasets, REALFP and SYNTHFP, for a novel task called Fermi Problems (FPs). This task necessitates the amalgamation of various reasoning abilities, including suitable abstractions, commonsense knowledge, and creative synthesis. Recently, mathematical reasoning has incorporated multiple input modalities such as textual, visual, and tabular data (Lu et al., 2023c; Zhu et al., 2021; Zhao et al., 2022). For instance, TabMWP (Lu et al., 2023c) presents a new dataset comprising 38,431 open-domain grade-level problems that necessitate mathematical reasoning across both textual and tabular data. Zhu et al. (Zhu et al., 2021) introduced a large-scale QA dataset, TAT-QA, encompassing both tabular and textual data. This dataset often requires numerical reasoning—such as addition, subtraction, multiplication, division, counting, comparison/sorting, and their combinations—based on real financial reports to infer answers. Additionally, MultiHiertt (Zhao et al., 2022) is a proposed dataset consisting of QA pairs derived from Multi Hierarchical Tabular and Textual data, where each document in this dataset contains multiple tables and longer unstructured texts.